Kyle has two more than three fourths as many nickels as dimes. Kyle has $8.35, How many of each coin does hehave?
Answer by stanbon(75887)

Kyle has two more than three fourths as many nickels as dimes. Kyle has $8.35, How many of each coin does hehave?
----
Equations:
d = (3/4)n +2
5n + 10d = 835 cents
=========================
Modify the 2nd equation:
n + 2d = 167
----
Substitute for "d" and solve for "n":
n + 2[(3/4)n + 2] = 167
n + (3/2)n + 4 = 167
(5/2)n = 163
n = (2/5)(163)
----
n is not a whole number.
Please check the statement of the problem.
--------------
Cheers,
Stan H.
===============